The length of a rectangle is 3 times it’s width. The perimeter of the rectangle is 72 meters. Find the dimensions
Answer by ikleyn(52781)

.

    x + 3x = 72/2 = 36


    4x            = 36

     x            = 36/4 = 9.


ANSWER.  The dimensions are 9 and 27 meters.

Solved.